# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ II.3 FT-IR
Figure 3: Sample plot from Section III.2.2 (see activity1.ipynb found in the Supporting Information). (A) Time-domain signal composed of three sine waves with frequency values of $\omega_{1}=4$ Hz, $\omega_{2}=6$ Hz, and $\omega_{3}=8$ Hz. (B) The frequency-domain signal obtained via a FT of the time-domain signal (Sum) showing distinct peaks at each frequency $\{\omega_{1},~{}\omega_{2},~{}\omega_{3}\}$. (c) Screenshot from a Jupyter notebook highlighting the interactive portion of Activity 1 (Sec. III.2.2) By adjusting the sliders for each frequency, the students can control the frequencies that create the musical chord (the black curve in panel A). Furthermore, students can change the duration of the chord by increasing/decreasing the number of sample points $N$ and modify the limits of the $x$-axis in panel (A) by entering different upper_xlim values.
@@
The critical difference between dispersive IR and FT-IR experiments is the Michelson interferometer. This interferometer allows one to probe a sample with all wavelengths of IR light simultaneously and obtain an interferogram, a temporal signal that encodes the interference of all light transmitted by the sample. The Michelson interferometer uses a beam splitter to split an incident IR light source into two beams: one reflects off a fixed mirror and the other reflects off a moving mirror. The moving mirror causes a path length difference between the two beams. Thus, the two beams interfere (constructively or destructively) when they recombine at the beam splitter. This light then passes through the sample, and the detector measures the total intensity of the transmitted IR wavelengths, yielding the desired interferogram. Similar to the FID in an NMR experiment, the interferogram contains the combined information of all transmitted wavelengths in the sample. Thus, one can employ the FT to disentangle this information. However, this spectrum will also contain information about the atmosphere surrounding the sample. One isolates the signal of the chemical sample by performing an additional FT-IR experiment of the atmosphere in the absence of the sample allowing us to obtain a background-subtracted spectrum.
To convert this interferogram into the conventional IR spectrum that shows percent transmittance (%$(\%T)$ as a function of frequency measured in wavenumbers (${\rm cm^{-1}}$), one converts the $x$-axis units from Hz to ${\rm cm^{-1}}$. One then obtains the absorption $A$ spectrum by applying the FT directly on the interferogram. Next, one employs $A=-\log(T)$ to obtain the transmittance ($T$), and computes the ratio of the transmittance of the sample $T_{\rm s}$ and the transmittance of the background $T_{\rm b}$, i.e., %$\%T=T_{\rm s}/T_{\rm b}$.
@@ III.1 Prelab
Students begin with a four-part prelab assignment (see prelab.pdf and prelab.ipynb in the Supporting Information) designed to prepare students with the math and Python concepts necessary for completing the activities. In Part 1, students focus on useful trigonometry and calculus concepts, including sinusoids, complex numbers, and integration, and employ these to understand the structure of the analytical form of the FT (Eq. (4)). Part 2 provides a tutorial for students to set up their Google Colab notebook. In Part 3, students use Google Colab to practice coding in Python. Finally, in Part 4 students are introduced to the basics of plotting in Python. We have designed these prelab activities to be modular, so that instructors can easily assign only the relevant material for their students, if, for instance, students have prior experience with Python.
III.2 Activity 1: Hearing the FT
@@
In Activity 1 (see activity1.ipynb in the Supporting Information), students begin developing their qualitative understanding of the FT with audio examples (musical chords generated in Python) and build a quantitative understanding by focusing on specific computations involving Eq. (4). This activity guides students to translate math into code, thereby employing a direct implementation of Eq. (4). This builds student confidence in coding and helps them observe the inefficiency of directly implementing Eq. (4), as discussed in Sec. II.1.3. Finally, they explore the practical implementation of the FT using the FFT algorithm, noting the significant increase in computational efficiency.
III.2.1 Part 1: Visualizing sinusoids
Learning goal: Students will gain familiarity with Python coding, produce high-quality figures in the matplotlib Python library, and apply their skills to visualizing time domain signals.
@@ III.2.1 Part 1: Visualizing sinusoids
III.2.2 Part 2: Listening for qualitative understanding
@@
In Part 2, students focus on qualitatively understanding the FT by analyzing how changes in the time-domain audio signals affect the frequency-domain spectrum. Specifically, students use interactive sliders to control the following parameters: frequency of each wave, length of the signal, and dampening factors. To reinforce sense-based learning, we have designed the activity such that students can listen to the audio generated with these parameters by pressing the ‘play’ button (Fig. 3C).
III.2.3 Part 3: Building quantitative understanding
Learning goal: Students will develop familiarity with Python programming by completing and analyzing pre-written code that directly translates the mathematical formulation of the FT in Eq. (3) to code. Students will observe how the FT works numerically by evaluating it at each frequency present in the signal and appreciate qualitatively that each peak in the frequency domain corresponds to a frequency present in the temporal signal. This exercise helps students begin to answer questions Q1 and Q2 from Sec. II.1.
In Part 3, the students begin exploring the quantitative aspects of the FT, while continuing to develop their qualitative understanding. First, we introduce the DFT (Eq. (6)) as the discrete analogue of Eq. (4). Then, the coding exercises guide students to compute the FT of the sinusoids shown in Fig. 3A. To minimize the complexity of the code while enabling students to understand the numerical components of the FT, students work with code to evaluate the FT (Eq. (6)) at a single frequency. Specifically, students evaluate the FT at each of the three frequencies, $\omega_{1}=4$ Hz, $\omega_{1}=6$ Hz, and $\omega_{1}=8$ Hz. Importantly, students will find that each computation results in a nonzero value because each of these frequencies are present in the input temporal signal. Students will additionally show that Eq. (6) is zero when evaluated at $\omega_{4}=10$ Hz since this frequency is not present in the temporal signal. Together, these computations demonstrate the origins of each peak in Fig. 3B.
III.2.4 Part 4: Observing the practical advantages of the FFT
@@
In Part 4, we introduce students to the practical (i.e., efficient) implementation of the FT while continuing to mine the FT for quantitative insight. We start by quantifying the idea of efficiency. Specifically, students compare the performance of their FT code to that of the FFT algorithm accessible through the numpy library. Students work with a built-in Python package (tqdm) to time their FT and FFT codes for datasets of increasing length. Although the remaining activities utilize the FFT, the intuition gained from the previous exercises is crucial to understanding the conceptual framework of the FT. The mathematics of the FFT are beyond the scope of these exercises and we refer the interested reader to Ref. Pres et al., 1992.
@@ III.3 Activity 2: Application to FT-NMR and FT-IR
Figure 4: Illustration of the FT in NMR and IR spectroscopy. (A) Model free induction decay (FID) given in Eq. (7) with $\omega_{1}=4$ Hz, $\omega_{2}=10$ Hz, and $T_{2}=0.5$ s. (B) Frequency spectrum obtained via the FFT of the FID in panel (A) showing distinct peaks centered at both frequencies $\{\omega_{1},\omega_{2}\}$ present in the FID. (C) Model interferogram of 3-pentanone constructed from sinuoids with frequencies corresponding to the following wavenumbers: $\omega_{1}=1200$ cm-1, $\omega_{2}=1720$ cm-1, $\omega_{3}=2900$ cm-1, $\omega_{4}=3000$ cm-1, $\omega_{5}=3500$ cm-1 (translucent curves). The background curve corresponds to a measurement made in the absence of a sample and thus all wavelengths are transmitted. The measurement curve corresponds to a measurement made in the presence of a sample where the sample absorbs 10% of light corresponding to $\omega_{2}=1720$ cm-1 and 5% of light corresponding to $\omega_{4}=3000$ cm-1. The inset magnifies the region enclosed in the red box. (D) Background subtracted transmittance ($T$) spectrum as a function of wavenumbers showing peaks at the frequencies absorbed by the chemical sample.
In Activity 2 (see activity2.ipynb in the Supporting Information), students bridge their qualitative and practical understanding of the FT with their experiences in the chemistry classroom. Specifically, in this activity students apply the concepts from Activity 1 to build and analyze FT-NMR and FT-IR datasets.
III.3.1 Part 1: Application to NMR spectroscopy
@@
${\rm FID}(t)=e^{-t/T_{2}}[\cos(\omega_{1}t)+\cos(\omega_{2}t)].$ (7)
We show this FID signal in Fig. 4A. Because this FID is aperiodic, students will apply the FT for functions defined on an infinite domain (Sec. II.1.2). The students apply the FFT algorithm to obtain the spectrum in Fig. 4B. As in Activity 1, students observe that each peak in the frequency domain is centered at one of the frequencies in Eq. (7). In addition to varying frequencies, students are asked to vary the amplitudes and phases of the sinusoids and the spin-spin relaxation constant, $T_{2}$, and to observe the resultant effects on the spectrum. Students will observe that: (i) changing amplitudes of the FID sinusoids results in a proportional change to the peak heights in the corresponding spectrum; (ii) increasing the phase from 00 to $2\pi$, the peaks change from slightly negative to completely positive; and (iii) decreasing $T_{2}$ causes peak widths to grow.
III.3.2 Part 2: Application to IR spectroscopy
Learning goal: Students will apply the FT to process and analyze a model IR interferogram. This activity will reiterate that peaks in the IR spectrum correspond to the wavelengths that transmit through the sample and onto the detector in an IR experiment. Students will review unit conversion by converting the intensity to % transmittance and frequency from Hertz (Hz) to wavenumbers in reciprocal centimeters (cm-1).
Here, students construct and analyze a fictitious FT-IR spectrum of 3-pentanone. We say “fictitious” because, for simplicity, we reduce the complexity of the light source (experimentally, a Globar) irradiating the sample, to an incident source composed of only 5 wavelengths: 1200 cm-1, 1720 cm-1, 2900 cm-1, 3000 cm-1, and 3500 cm-1. We choose these frequencies to illustrate two frequencies where one expects absorption and three frequencies where no absorption occurs. Students generate the corresponding sinusoids for each wavelength, shown as translucent curves in Fig. 4C. In the foreground of Fig. 4C, we illustrate the ‘background’ measurement where one measures all of the incident light at the detector due to the absence of a sample. We then show a ‘measured’ interferogram curve constructed under two assumptions: (i) that our sample absorbs 10% of the 1720 cm-1 light, 5% of the 3000 cm-1 light, and (ii) that all other wavelengths of light pass through the sample. Students combine their acquired knowledge of the FT from Sec. II.1 and FT-IR theory in Sec. II.3 to compute the spectrum of percent transmittance in Fig. 4D. Finally, we ask students to consider the differences between this modeled spectrum and an experimental spectrum from Ref. NIST Chemistry WebBook, 1977, encouraging them to reflect on the coupling of normal modes within a molecule.
@@ III.4 Activity 3: Processing your own FID
In Activity 3 (see activity3.ipynb in the Supporting Information), students synthesize the skills developed in Secs. III.2 and III.3 to process an FID signal corresponding to an experimental FT-NMR signal for 2-butanone in deuterated chloroform solvent (CDCl3), collected on a Varian 500 MHz NMR spectrometer. We ask students to plot and manipulate the raw FID file obtained from the FT-NMR experiment. Specifically, we first ask students to subtract the average of the FID measurements to ensure that the FID decays to zero at long times, satisfying the FT requirement that ${\rm FID}(t)\rightarrow 0$ as $t\rightarrow\infty$ (Sec. II.1). We then provide code that zero-pads the spectrum and explain that this manipulation ensures that the length of the FID is a power of 2, maximizing the efficiency of the FFT (Section II.1). Finally, we introduce students to the Lanczos and Hanning windowing functions that improve the signal-to-noise (S/N) ratio.Prabhu (2018) We invite students to explore the effects of each windowing function as well as the combination or absence of both windowing functions. After post-processing the spectrum, we ask students to compare the processed (windowed) spectrum to the unprocessed spectrum (without windowing). Finally, we provide students with the chemical formula for 2-butanone and ask them to identify the molecule and label the corresponding signals on their spectrum. In Fig. 5, we display the resulting spectrum that students are expected to generate and compare it to the spectrum we have processed using Mnova,Mestrelab Research (2024) a widely used FT-NMR post-processing software. This activity aims to empower students with the tools to process and analyze their own FID datasets.
@@ III.5 Implementation of this lab activity
At Trinity, the majority of students had no prior programming experience and were not previously familiar with the mathematics of the FT. At Chatham, students had been doing weekly exercises that included some programming for a few weeks, but for almost all of them, this was their first introduction to the FT. We administered the prelab exercise during class time for students at Trinity, who did not have prior coding experience. All students completed the prelab independently without assistance from the instructor, taking $\sim$30 minutes to complete the math portion (Part 1) and $\sim$45 minutes to complete the coding portion (Part 2-4). At Chatham and Trinity, students completed the three activities over two successive three-hour lab periods. Each student completed the activities independently. Activity 1 took students $\sim$90 minutes, Activity 2 required $\sim$100 minutes, and Activity 3 required $\sim$120 minutes.
@@ IV Suggested Modifications
We have designed these activities to be modular so that instructors can choose the sections that fit their curriculum or time constraints. For example, if students are already familiar with Python programming, instructors can omit sections that cover the introduction to Python. We suggest three modifications to reduce the time required for the students to complete these activities. First, instructors can adopt a shorter version of the laboratory by removing either Part 1 (Sec. III.3.1) or Part 2 (Sec. III.3.2) from Activity 2. This reduces the total time by $\sim$50 minutes. Second, the instructor could remove Activity 3, reducing the total time by $\sim$120 minutes. Third, we have provided a streamlined version of the activities that combines Activities 1 and 2 but does not include as many plotting manipulations with matplotlib. Additionally, instructors can expand on Activity 3 by comparing the data collected on a higher resolution instrument (e.g., in Activity 3, we used 500 MHz) to a lower resolution instrument (e.g., a benchtop NMR with 90 MHz). Lastly, the topics serve as an introduction to the Fourier analysis toolkit which instructors can expand on in many ways, such as with deeper discussions of the FFT algorithm, zero-padding, and windowing functions. Furthermore, this sets the framework for introducing more complex topics such as aliasing, noise reduction, and phase correction.

### v2

@@ II.3 FT-IR
[…]
@@ III.1 Prelab
Students begin with a four-part prelab assignment (see Prelab-Questions.pdf and Prelab.ipynb in the Supporting Information) designed to prepare students with the math and Python concepts necessary for completing the activities. In Prelab 1.1, students focus on useful trigonometry and calculus concepts, including sinusoids, complex numbers, and integration, and employ these to understand the structure of the analytical form of the FT (Eq. (4)). Prelab 1.2 provides a tutorial for students to set up their Google Colab notebook. In Prelab 1.3, students use Google Colab to practice coding in Python. Finally, in Prelab 1.4 students are introduced to the basics of plotting in Python. We have designed these prelab activities to be modular, so that instructors can easily assign only the relevant material for their students, if, for instance, students have prior experience with Python.
Figure 3: Sample plots from Activities 1.2-1.3 (See Secs. III.2.2 & III.2.3). (A) Screenshot from a Jupyter notebook highlighting the interactive portion of Activity 1.2. By adjusting the sliders for each frequency, students can control the frequencies that create the musical chord (the black curve in panel B). Students can also change the duration of the chord by increasing/decreasing the number of sample points $N$ and modify the limits of the $x$-axis in panel (B) by entering different upper_xlim values. (B) Time-domain signal composed of three sine waves with frequency values of $\omega_{1}=4$ s-1, $\omega_{2}=6$ s-1, and $\omega_{3}=8$ s-1. Students can set each of these values using the sliders in panel (A). (C) The frequency-domain signal obtained via a FT of the time-domain signal in panel (B), showing distinct peaks at each frequency $\{\omega_{1},~{}\omega_{2},~{}\omega_{3}\}$.
III.2 Activity 1: Applying the FT to audio signals
@@
In Activity 1 (see Activity_1.ipynb in the Supporting Information), students begin developing their qualitative understanding of the FT with audio examples by determining the frequencies and amplitudes of the pure tones that contribute to musical chords generated in Python. Students then build a quantitative understanding by focusing on specific computations involving Eq. (4). This activity guides students to translate math into code, thereby employing a direct implementation of Eq. (4). This builds student confidence in coding and helps them observe the inefficiency of directly implementing Eq. (4), as discussed in Sec. II.1.3. Finally, they explore the practical implementation of the FT using the FFT algorithm, noting the significant increase in computational efficiency.
III.2.1 Activity 1.1: Visualizing sinusoids
Learning goal: Students will gain familiarity with Python coding, produce high-quality figures in the matplotlib Python library, and apply their skills to visualize time domain signals.
@@
III.2.2 Activity 1.2: Listening for qualitative understanding
@@
In Activity 1.2, students focus on qualitatively understanding the FT by analyzing how changes in the time-domain audio signals affect the frequency-domain spectrum. Specifically, students use interactive sliders to control the following parameters: frequency of each wave, length of the signal, and dampening factors. To reinforce sense-based learning, we have designed the activity such that students can listen to the audio generated with these parameters by pressing the ‘play’ button (Fig. 3A).
III.2.3 Activity 1.3: Building quantitative understanding
Learning goal: Students will develop familiarity with Python programming by completing and analyzing pre-written code that directly translates the mathematical formulation of the FT in Eq. (3) to code. Students will observe how the FT works numerically by evaluating it at each frequency present in a signal and appreciate qualitatively that each peak in the frequency domain corresponds to a frequency present in the temporal signal. This exercise helps students begin to answer questions Q1 and Q2 from Sec. II.1.
In Activity 1.3, the students begin exploring the quantitative aspects of the FT, while continuing to develop their qualitative understanding. First, we introduce the DFT (Eq. (6)) as the discrete analogue of Eq. (4). Then, the coding exercises guide students to compute the FT of the sinusoids shown in Fig. 3B. To minimize the complexity of the code while enabling students to understand the numerical components of the FT, students work with code to evaluate the FT (Eq. (6)) at a single frequency. Specifically, students evaluate the FT at each of the three frequencies, $\omega_{1}=4$ s-1, $\omega_{1}=6$ s-1, and $\omega_{1}=8$ s-1. Importantly, students will find that each computation results in a nonzero value because each of these frequencies are present in the input temporal signal. Students will additionally show that Eq. (6) is zero when evaluated at $\omega_{4}=10$ s-1 since this frequency is not present in the temporal signal. Together, these computations demonstrate the origins of each peak in Fig. 3C.
III.2.4 Activity 1.4: Practical advantages of the FFT
@@
In Activity 1.4, we introduce students to the practical (i.e., efficient) implementation of the FT while continuing to mine the FT for quantitative insight. We start by quantifying the idea of efficiency. Specifically, students compare the performance of their FT code to that of the FFT algorithm accessible through the numpy library. Students work with a built-in Python package (tqdm) to time their FT and FFT codes for datasets of increasing length. Although the remaining activities utilize the FFT, the intuition gained from the previous exercises is crucial to understanding the conceptual framework of the FT.
@@ III.3 Activity 2: Application to FT-NMR and FT-IR
[…]
In Activity 2 (see Activity_2.ipynb in the Supporting Information), students bridge their qualitative and practical understanding of the FT with their experiences in the chemistry classroom. Specifically, in this activity students apply the concepts from Activity 1 to build and analyze FT-NMR and FT-IR datasets.
III.3.1 Activity 2.1: Application to NMR spectroscopy
@@
${\rm FID}(t)=e^{\tfrac{-t}{T_{2}}}[A_{1}\cos(\omega_{1}t+\phi_{1})+A_{2}\cos(\omega_{2}t+\phi_{2})],$ (7)
where $\{A_{1},A_{2}\}$, $\{\omega_{1},\omega_{2}\}$, and $\{\phi_{1},\phi_{2}\}$ are the amplitudes, angular frequencies, and phases of each sinusoid. We show this FID signal with $A_{1}=A_{2}=1$, $\omega_{1}=4$ s-1, $\omega_{2}=10$ s-1, and $\phi_{1}=\phi_{2}=0$ in Fig. 4A. Because this FID is aperiodic, students will apply the FT as we discussed in Sec. II.1.2. The students apply the FFT algorithm to obtain the spectrum in Fig. 4B. As in Activity 1, students observe that each peak in the spectrum (Fig. 4B) is centered at one of the frequencies in Eq. (7). In addition to varying frequencies, students are asked to vary the amplitudes and phases of the sinusoids and the spin-spin relaxation time, $T_{2}$, and to observe the resultant effects on the spectrum. Students will observe that: (i) changing amplitudes of the FID sinusoids results in a proportional change to the peak heights in the corresponding spectrum; (ii) increasing the phase from 00 to $2\pi$, the peaks change from slightly negative to completely positive; and (iii) decreasing $T_{2}$ causes peak widths to grow.
III.3.2 Activity 2.2: Application to IR spectroscopy
Learning goal: Students will apply the FT to process and analyze a model IR interferogram. This activity will reiterate that peaks in the IR spectrum correspond to the frequencies that transmit through the sample and onto the detector in an IR experiment.
[…]
@@ III.4 Activity 3: Processing your own FID
In Activity 3 (see Activity_3.ipynb in the Supporting Information), students synthesize the skills developed in Secs. III.2 and III.3 to process an FID signal corresponding to an experimental FT-NMR signal for 2-butanone in deuterated chloroform solvent (CDCl3), collected on a Varian 500 MHz NMR spectrometer. We ask students to plot and manipulate the raw FID file obtained from the FT-NMR experiment. Specifically, we first ask students to subtract the average of the FID measurements to ensure that the FID decays to zero at long times, satisfying the FT requirement that ${\rm FID}(t)\rightarrow 0$ as $t\rightarrow\infty$ (Sec. II.1). We then provide code to zero-pad the end of the temporal signal (i.e., append zeros), and explain that this allows one to adjust the resolution of spectrum as well as ensuring that the length of the FID is a power of 2, maximizing the efficiency of the FFT (Sec. II.1). Finally, we introduce students to the Lanczos and Hanning windowing functions that improve the signal-to-noise (S/N) ratio.Prabhu (2018) We invite students to explore the effects of each windowing function as well as the combination or absence of both windowing functions. After post-processing the spectrum, we ask students to compare the processed (windowed) spectrum to the unprocessed spectrum (without windowing). Finally, we provide students with the chemical formula for 2-butanone and ask them to identify the molecule and label the corresponding signals on their spectrum. In Fig. 5, we display the resulting spectrum that students are expected to generate and compare it to the spectrum we have processed using Mnova,Mestrelab Research (2024) a widely used FT-NMR post-processing software. This activity aims to empower students with the tools to process and analyze their own FID datasets.
@@ III.5 Implementation of this lab activity
At Trinity, the majority of students had no prior programming experience and were not previously familiar with the mathematics of the FT. At Chatham, students had been doing weekly exercises that included some programming for a few weeks, but for almost all of them, this was their first introduction to the FT. We administered the prelab exercise during class time for students at Trinity, who did not have prior coding experience. All students completed the prelab independently without assistance from the instructor, taking $\sim$30 minutes to complete the math portion (Prelab 1.1) and $\sim$45 minutes to complete the coding portion (Prelab 1.2-1.4). At Chatham and Trinity, students completed the three activities over two successive three-hour lab periods. Each student completed the activities independently. Activity 1 took students $\sim$90 minutes, Activity 2 required $\sim$100 minutes, and Activity 3 required $\sim$120 minutes.
@@ IV Suggested Modifications
We have designed these activities to be modular so that instructors can choose the sections that fit their curriculum or time constraints. For example, if students are already familiar with Python programming, instructors can omit sections that cover the introduction to Python.
As designed, the three activities take approximately two 3-hour periods and fit well into an upper-division chemistry laboratory course. We recognize that this is a significant time commitment, and encourage instructors to adapt the modules as needed to fit their time constraints. Here, we suggest three specific modifications that reduce the time required for the students to complete these activities. First, instructors can adopt a shorter version of the laboratory by removing either Activity 2.1 or Activity 2.2. Removing Activity 2.2, frames the activities solely with a focus on NMR spectroscopy as an analytical technique. Removing Activity 2.1 or Activity 2.2 reduces the total activity time by $\sim$50 minutes.
Second, for instructors with only one $\sim$3 hour lab period available for this activity, we suggest omitting Activity 3 Activity 3, which reduces the total activity time by $\sim$120 minutes. With this modification, students still investigate the role of the FT in both NMR and IR spectroscopy in Activity 2, but without the additional data processing considerations for “real” experimental NMR data given in Activity 3.
Third, instructors in lecture courses are also likely to find the content covered in this activity of interest. For this purpose, we have provided a streamlined version of the activities combining Activities 1 and 2 that is appropriate for a $\sim$1-hour class period. The interested reader can find this streamlined activity (labeled Streamlined.ipynb) in the Supporting Information. In this version, students work through the same conceptual material, but with more guidance in the Python coding and less emphasis on plotting with the matplotlib library. This further reduces any barriers due to lack of familiarity with Python while maintaining the ability to meet the learning objectives of the activity in a condensed format.
For interested instructors or students, there are many possibilities for extending these activities, especially Activity 3. If instructors want to add an experimental component to this activity, students could collect their own experimental NMR data on their molecule of choice, extract the FID, and use that data to conduct Activity 3. This would also lend itself to an exploration of the impact of instrument resolution on the collected FID and resultant spectrum. For example, instructors could expand on Activity 3 by comparing the data collected on a higher resolution instrument (e.g., in Activity 3, we used 500 MHz) to a lower resolution instrument (e.g., a benchtop NMR with 90 MHz). The topics included in these activities serve as an introduction to the Fourier analysis toolkit, on which instructors can expand on in many ways, such as with deeper discussions of the FFT algorithm, zero-padding, and windowing functions. Furthermore, this sets the framework for introducing more complex topics such as aliasing, noise reduction, and phase correction.